At 30ºC,Kp for the dissociation reaction SO2Cl2(g)⇌SO2(g)+Cl2(g) is 2.9×10^−2 atm. If the total pressure is 1 atm, the degree of dissociation of SO2Cl2 is (assume, 1−α^2=1): (A) 83% (B) 13% (C) 17% (D) 29%

Grade:11

1 Answers

Prasanna shetty
4 years ago
                        SO2Cl2 = SO2 + Cl2
                             1          0         0
At equilibrium        1-x       x         x
mole fraction         1+x    1+x     1+x
 
         kp = SO2 X  Cl2
                   SO2Cl2
         kp =   x/1+x .P   x/1+x. P
                         1-x/1+x .P
After simplication,
        kp =  x2
              1 – x2        ( Here 1 – x2 =1, it is given), 
 
so equatiion turns into kp = x2
 
substitute the value of kp= 2.9 x 10-2 , then x = 1.7 X 10-1   or 17%.